OpenStudy (lxdrpr):

Write an equation for the line that is parallel to the given line and that passes through the given point. y = 2x+ 7; (3, 11) how would you solve this?? how would the y intercept change and how does the point affect this?

OpenStudy (shadowlegendx):

Do you know what a parallel line is?

OpenStudy (lxdrpr):

yes so far i have y=2x but idk how the y intercept would change

OpenStudy (lxdrpr):

the slope remains the same ik that

OpenStudy (shadowlegendx):

Yes, so we know that the slope of this new line will be 2x, but we also know that it must pass through (3,11). So we must determine the y intercept. This requires a little, head work. Our line rises by 2 units and runs by 1 unit 2/1 So wherever it intercepts along the y axis, it would have to eventually reach (3,11) I can determine the y intercept, by bringing subtracting 1 unit from (3,11) for the run, and 2 units from 11 for the rise. I get (2,9). Repeat, (1,7). Repeat (0, 5)

OpenStudy (shadowlegendx):



OpenStudy (lxdrpr):

ok so you remove the units until the y intercept is the one found so it would be y=2x+5

OpenStudy (shadowlegendx):

Yes, as it's parallel to y = 2x + 7 and will eventually reach (3,11) on the graph

OpenStudy (lxdrpr):

ok thank you!

OpenStudy (shadowlegendx):

No problem

OpenStudy (shadowlegendx):

My way is a little unorthodox, so just so you have the formal way to refer to... Point Slope Formula: \[y - y _{1} = m(x - x _{1})\] Where, m = slope x1 and y1 = any point In this problem, m = 2x Point = (3,11) Input \[y - 11 = 2(x - 3)\] Distribute 2 \[y - 11 = 2x - 6\] Add 11 to both sides \[y = 2x + 5\]